OpenStudy (anonymous):

A telephone company charges customers a certain amount for the first minute of a long distance call and another amount for each additional minute. A customer makes two calls to the same city—a 31-minute call for $1.94 and a 13-minute call for $0.86. If there is a federal tax rate of 3.7% and a state tax rate of 7.4% on all long distance calls, find, to the nearest minute, the longest call to the same city whose cost will not exceed $9.00.

OpenStudy (anonymous):

for the first half of the problem i got .14 for the cost of the first minute and then .06 for the additional minutes.. i need help solving the last question about the federal tax

OpenStudy (anonymous):

I'd say add the taxes together 3.7+7.4 = 11.1% then the equation becomes y = (11.1/100)*( 0.14 + 0.6(m-1) ) does that work?

OpenStudy (anonymous):

am i soppose to solve for m? or y?

OpenStudy (anonymous):

im confused and its .06

OpenStudy (anonymous):

you solve for m (mintute) given that y (money) equals $9

OpenStudy (anonymous):

@xon1300 In your formula, \(y\) would only be the tax. It wouldn't include the cost of the call.

OpenStudy (anonymous):

so \[y = (111.1/100)*( 0.14 + 0.06(m-1) )\] when y=9 m= what?

OpenStudy (anonymous):

There, that one's better. :)

OpenStudy (anonymous):

hard to count ones, they marge together, lol

OpenStudy (anonymous):

Haha yeah!

OpenStudy (anonymous):

Thank you so much!! thats the same thing i did and thats why i couldn't figure it out i hade 11.1